Find the common factors of the given terms.
16 x^3 , - 4x^2 , 32 x

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Factors of 16x^3, -4x^2and 32x
16 x^3= 2 × 2 × 2 × 2 × x × x × x
- 4x^2= -1 × 2 × 2 × x × x
32x = 2 × 2 × 2 × 2 × 2 × x
Common factors of 16 x^3, - 4x^2and 32x are 2,2, x and, 2 × 2 × x = 4x

Thanks